# Euclid: Early Release Observations – The star cluster systems of the Local Group dwarf galaxies IC 10 and NGC 6822††thanks: This paper is published on behalf of the Euclid Consortium

### v1

@@ Acknowledgements.
- Adamo et al. (2024) Adamo, A., Bradley, L. D., Vanzella, E., et al. 2024, Nature, 632, 513
- Adamo et al. (2015) Adamo, A., Kruijssen, J. M. D., Bastian, N., Silva-Villa, E., & Ryon, J. 2015, MNRAS, 452, 246
- Adamo et al. (2017) Adamo, A., Ryon, J. E., Messa, M., et al. 2017, ApJ, 841, 131
- Astropy Collaboration et al. (2022) Astropy Collaboration, Price-Whelan, A. M., Lim, P. L., et al. 2022, ApJ, 935, 167
- Barmby et al. (2009) Barmby, P., Perina, S., Bellazzini, M., et al. 2009, AJ, 138, 1667
- Barnard (1884) Barnard, E. E. 1884, Astronomische Nachrichten, 110, 125
- Bastian et al. (2012) Bastian, N., Adamo, A., Gieles, M., et al. 2012, MNRAS, 419, 2606
- Battaglia et al. (2022) Battaglia, G., Taibi, S., Thomas, G. F., & Fritz, T. K. 2022, A&A, 657, A54
- Battinelli et al. (2006) Battinelli, P., Demers, S., & Kunkel, W. E. 2006, A&A, 451, 99
- Baumgardt et al. (2020) Baumgardt, H., Sollima, A., & Hilker, M. 2020, PASA, 37, e046
- Bekki & Freeman (2003) Bekki, K. & Freeman, K. C. 2003, MNRAS, 346, L11
- Bennet et al. (2024) Bennet, P., Patel, E., Sohn, S. T., et al. 2024, ApJ, 971, 98
- Bianchini et al. (2014) Bianchini, P., Renaud, F., Gieles, M., & Varri, A. L. 2014, MNRAS, 447, L40
- Bica et al. (2020) Bica, E., Westera, P., Kerber, L. d. O., et al. 2020, AJ, 159, 82
- Bradley et al. (2023) Bradley, L., Sipőcz, B., Robitaille, T., et al. 2023, astropy/photutils: 1.8.0
- Bressan et al. (2012) Bressan, A., Marigo, P., Girardi, L., et al. 2012, MNRAS, 427, 127
- Brodie & Strader (2006) Brodie, J. P. & Strader, J. 2006, ARA&A, 44, 193
- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
- Buchner et al. (2014) Buchner, J., Georgakakis, A., Nandra, K., et al. 2014, A&A, 564, A125
- Cardelli et al. (1989) Cardelli, J. A., Clayton, G. C., & Mathis, J. S. 1989, ApJ, 345, 245
- Carlsten et al. (2022) Carlsten, S. G., Greene, J. E., Beaton, R. L., & Greco, J. P. 2022, ApJ, 927, 44
- Carnall et al. (2018) Carnall, A. C., McLure, R. J., Dunlop, J. S., & Davé, R. 2018, MNRAS, 480, 4379
- Cerny et al. (2023) Cerny, W., Drlica-Wagner, A., Li, T. S., et al. 2023, ApJ, 953, L21
- Cerny et al. (2021) Cerny, W., Pace, A. B., Drlica-Wagner, A., et al. 2021, ApJ, 910, 18
- Chandar et al. (2000) Chandar, R., Bianchi, L., & Ford, H. C. 2000, AJ, 120, 3088
- Chandar et al. (2005) Chandar, R., Leitherer, C., Tremonti, C. A., et al. 2005, ApJ, 628, 210
- Chandar et al. (2016) Chandar, R., Whitmore, B. C., Dinino, D., et al. 2016, ApJ, 824, 71
- Chandar et al. (2010) Chandar, R., Whitmore, B. C., Kim, H., et al. 2010, ApJ, 719, 966
- Chen et al. (2015) Chen, Y., Bressan, A., Girardi, L., et al. 2015, MNRAS, 452, 1068
- Cole et al. (2017) Cole, A. A., Weisz, D. R., Skillman, E. D., et al. 2017, ApJ, 837, 54
- Cook et al. (2019) Cook, D. O., Lee, J. C., Adamo, A., et al. 2019, MNRAS, 484, 4897
- Côté et al. (1998) Côté, P., Marzke, R. O., & West, M. J. 1998, ApJ, 501, 554
- Crnojević et al. (2016) Crnojević, D., Sand, D. J., Zaritsky, D., et al. 2016, ApJ, 824, L14
- Cuillandre et al. (2025) Cuillandre, J. C., Bertin, E., Bolzonella, M., et al. 2025, A&A, 697, A6
- de Blok & Walter (2000) de Blok, W. J. G. & Walter, F. 2000, ApJ, 537, L95
- Deason et al. (2014) Deason, A., Wetzel, A., & Garrison-Kimmel, S. 2014, ApJ, 794, 115
- Deger et al. (2022) Deger, S., Lee, J. C., Whitmore, B. C., et al. 2022, MNRAS, 510, 32
- Dell’Agli et al. (2018) Dell’Agli, F., Di Criscienzo, M., Ventura, P., et al. 2018, MNRAS, 479, 5035
- Demers et al. (2006) Demers, S., Battinelli, P., & Kunkel, W. E. 2006, ApJ, 636, L85
- Demers et al. (2004) Demers, S., Battinelli, P., & Letarte, B. 2004, A&A, 424, 125
- Dotter et al. (2007) Dotter, A., Chaboyer, B., Jevremović, D., et al. 2007, AJ, 134, 376
- Eadie et al. (2022) Eadie, G. M., Harris, W. E., & Springford, A. 2022, ApJ, 926, 162
- Efremova et al. (2011) Efremova, B. V., Bianchi, L., Thilker, D. A., et al. 2011, ApJ, 730, 88
- Elmegreen & Efremov (1997) Elmegreen, B. G. & Efremov, Y. N. 1997, ApJ, 480, 235
- Elson et al. (1987) Elson, R. A. W., Fall, S. M., & Freeman, K. C. 1987, ApJ, 323, 54
- Euclid Collaboration: Cropper et al. (2025) Euclid Collaboration: Cropper, M., Al-Bahlawan, A., Amiaux, J., et al. 2025, A&A, 697, A2
- Euclid Collaboration: Jahnke et al. (2025) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2025, A&A, 697, A3
- Euclid Collaboration: Mellier et al. (2025) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2025, A&A, 697, A1
- Euclid Collaboration: Scaramella et al. (2022) Euclid Collaboration: Scaramella, R., Amiaux, J., Mellier, Y., et al. 2022, A&A, 662, A112
- Euclid Early Release Observations (2024) Euclid Early Release Observations. 2024, https://doi.org/10.57780/esa-qmocze3
- Fall & Zhang (2001) Fall, S. M. & Zhang, Q. 2001, ApJ, 561, 751
- Feroz & Hobson (2008) Feroz, F. & Hobson, M. P. 2008, MNRAS, 384, 449
- Feroz et al. (2009) Feroz, F., Hobson, M. P., & Bridges, M. 2009, MNRAS, 398, 1601
- Floyd et al. (2024) Floyd, M., Chandar, R., Whitmore, B. C., et al. 2024, AJ, 167, 95
- Forbes et al. (2018a) Forbes, D. A., Bastian, N., Gieles, M., et al. 2018a, Proceedings of the Royal Society of London Series A, 474, 20170616
- Forbes et al. (2018b) Forbes, D. A., Read, J. I., Gieles, M., & Collins, M. L. M. 2018b, MNRAS, 481, 5592
- Foster et al. (2014) Foster, C., Lux, H., Romanowsky, A. J., et al. 2014, MNRAS, 442, 3544
- Fouesneau & Lançon (2010) Fouesneau, M. & Lançon, A. 2010, A&A, 521, A22
- Fusco et al. (2012) Fusco, F., Buonanno, R., Bono, G., et al. 2012, A&A, 548, A129
- Fusco et al. (2014) Fusco, F., Buonanno, R., Hidalgo, S. L., et al. 2014, A&A, 572, A26
- Gallart et al. (1996) Gallart, C., Aparicio, A., & Vilchez, J. M. 1996, AJ, 112, 1928
- Garnett (1990) Garnett, D. R. 1990, ApJ, 363, 142
- Gatto et al. (2021) Gatto, M., Ripepi, V., Bellazzini, M., et al. 2021, MNRAS, 507, 3312
- Georgiev et al. (2010) Georgiev, I. Y., Puzia, T. H., Goudfrooij, P., & Hilker, M. 2010, MNRAS, 406, 1967
- Georgiev et al. (2009) Georgiev, I. Y., Puzia, T. H., Hilker, M., & Goudfrooij, P. 2009, MNRAS, 392, 879
- Gerbrandt et al. (2015) Gerbrandt, S. A. N., McConnachie, A. W., & Irwin, M. 2015, MNRAS, 454, 1000
- Gieren et al. (2006) Gieren, W., Pietrzyński, G., Nalewajko, K., et al. 2006, ApJ, 647, 1056
- Glatt et al. (2008) Glatt, K., Grebel, E. K., Sabbi, E., et al. 2008, AJ, 136, 1703
- Gouliermis et al. (2010) Gouliermis, D. A., Schmeja, S., Klessen, R. S., de Blok, W. J. G., & Walter, F. 2010, ApJ, 725, 1717
- Grasha et al. (2019) Grasha, K., Calzetti, D., Adamo, A., et al. 2019, MNRAS, 483, 4707
- Gratton et al. (2019) Gratton, R., Bragaglia, A., Carretta, E., et al. 2019, A&ARv, 27, 8
- Green (2018) Green, G. 2018, The Journal of Open Source Software, 3, 695
- Guy et al. (2022) Guy, L. P., Cuillandre, J.-C., Bachelet, E., et al. 2022, in Zenodo id. 5836022, Vol. 58, 5836022
- Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357
- Harris (1996) Harris, W. E. 1996, AJ, 112, 1487
- Harris & van den Bergh (1981) Harris, W. E. & van den Bergh, S. 1981, AJ, 86, 1627
- Higgs et al. (2021) Higgs, C. R., McConnachie, A. W., Annau, N., et al. 2021, MNRAS, 503, 176
- Hirschauer et al. (2020) Hirschauer, A. S., Gray, L., Meixner, M., et al. 2020, ApJ, 892, 91
- Hodge et al. (1988) Hodge, P., Kennicutt, Jr., R. C., & Lee, M. G. 1988, PASP, 100, 917
- Hodge (1977) Hodge, P. W. 1977, ApJS, 33, 69
- Hodge (1980) Hodge, P. W. 1980, ApJ, 241, 125
- Hubble (1925) Hubble, E. P. 1925, ApJ, 62, 409
- Hunt et al. (2025) Hunt, L. K., Annibali, F., Cuillandre, J. C., et al. 2025, A&A, 697, A9
- Hunter (2001) Hunter, D. A. 2001, ApJ, 559, 225
- Hurley & Mackey (2010) Hurley, J. R. & Mackey, A. D. 2010, MNRAS, 408, 2353
- Huxor et al. (2013) Huxor, A. P., Ferguson, A. M. N., Veljanoski, J., Mackey, A. D., & Tanvir, N. R. 2013, MNRAS, 429, 1039
- Huxor et al. (2014) Huxor, A. P., Mackey, A. D., Ferguson, A. M. N., et al. 2014, MNRAS, 442, 2165
- Huxor et al. (2008) Huxor, A. P., Tanvir, N. R., Ferguson, A. M. N., et al. 2008, MNRAS, 385, 1989
- Huxor et al. (2005) Huxor, A. P., Tanvir, N. R., Irwin, M. J., et al. 2005, MNRAS, 360, 1007
- Hwang et al. (2011) Hwang, N., Lee, M. G., Lee, J. C., et al. 2011, ApJ, 738, 58
- Hwang et al. (2014) Hwang, N., Park, H. S., Lee, M. G., et al. 2014, ApJ, 783, 49
- Jang et al. (2012) Jang, I. S., Lim, S., Park, H. S., & Lee, M. G. 2012, ApJ, 751, L19
- Jarrett et al. (2003) Jarrett, T. H., Chester, T., Cutri, R., Schneider, S. E., & Huchra, J. P. 2003, AJ, 125, 525
- Johnson et al. (2017) Johnson, L. C., Seth, A. C., Dalcanton, J. J., et al. 2017, ApJ, 839, 78
- Johnson et al. (2015) Johnson, L. C., Seth, A. C., Dalcanton, J. J., et al. 2015, ApJ, 802, 127
- Johnson et al. (2022) Johnson, L. C., Wainer, T. M., Torresvillanueva, E. E., et al. 2022, ApJ, 938, 81
- Jones et al. (2023) Jones, M. G., Karunakaran, A., Bennet, P., et al. 2023, ApJ, 942, L5
- Karachentsev & Tikhonov (1993) Karachentsev, I. D. & Tikhonov, N. A. 1993, A&AS, 100, 227
- Karim et al. (2024) Karim, N., Collins, M. L. M., Forbes, D. A., & Read, J. I. 2024, MNRAS, 530, 4936
- Khatamsaz et al. (2024) Khatamsaz, F., Abdollahi, M., Abdollahi, H., Javadi, A., & van Loon, J. T. 2024, arXiv e-prints, arXiv:2412.05646
- Kim et al. (2009) Kim, M., Kim, E., Hwang, N., et al. 2009, ApJ, 703, 816
- King (1966) King, I. R. 1966, AJ, 71, 64
- Kirby et al. (2013) Kirby, E. N., Cohen, J. G., Guhathakurta, P., et al. 2013, ApJ, 779, 102
- Krienke & Hodge (2004) Krienke, K. & Hodge, P. 2004, PASP, 116, 497
- Kroupa (2001) Kroupa, P. 2001, MNRAS, 322, 231
- Kruijssen (2019) Kruijssen, J. M. D. 2019, MNRAS, 486, L20
- Krumholz et al. (2019) Krumholz, M. R., McKee, C. F., & Bland-Hawthorn, J. 2019, MNRAS, 57, 227
- Lahén et al. (2025) Lahén, N., Rantala, A., Naab, T., et al. 2025, MNRAS
- Lang et al. (2010) Lang, D., Hogg, D. W., Mierle, K., Blanton, M., & Roweis, S. 2010, AJ, 139, 1782
- Larsen (2004) Larsen, S. S. 2004, A&A, 416, 537
- Larsen (2009) Larsen, S. S. 2009, A&A, 494, 539
[…]

### v3

@@ Acknowledgements.
- Probing the role of the galactic environment in the formation of stellar clusters, using M83 as a test bench. MNRAS 452 (1), pp. 246–260. External Links: Document, 1505.07475 Cited by: §5.1, §7.2.
- Legacy ExtraGalactic UV Survey with The Hubble Space Telescope: Stellar Cluster Catalogs and First Insights Into Cluster Formation and Evolution in NGC 628. ApJ 841 (2), pp. 131. External Links: Document, 1705.01588 Cited by: §1.
- Bound star clusters observed in a lensed galaxy 460 Myr after the Big Bang. Nature 632 (8025), pp. 513–516. External Links: Document, 2401.03224 Cited by: §1, item 2.
- The Astropy Project: Sustaining and Growing a Community-oriented Open-source Project and the Latest Major Release (v5.0) of the Core Package. ApJ 935 (2), pp. 167. External Links: Document, 2206.14220 Cited by: Data Availability.
- A Hubble Space Telescope/WFPC2 Survey of Bright Young Clusters in M31. III. Structural Parameters. AJ 138 (6), pp. 1667–1680. External Links: Document, 0909.4053 Cited by: §7.4.
- New Nebula near General Catalogue No 4510. Astronomische Nachrichten 110, pp. 125. External Links: Document Cited by: §2.1.
- Stellar clusters in M83: formation, evolution, disruption and the influence of the environment. MNRAS 419 (3), pp. 2606–2622. External Links: Document, 1109.6015 Cited by: §7.4.
- Gaia early DR3 systemic motions of Local Group dwarf galaxies and orbital properties with a massive Large Magellanic Cloud. A&A 657, pp. A54. External Links: Document, 2106.08819 Cited by: §2.1.
- Photometric survey of the polar ring galaxy NGC 6822. A&A 451 (1), pp. 99–108. External Links: Document, astro-ph/0603558 Cited by: §2.1.
- Formation of ω centauri from an ancient nucleated dwarf galaxy in the young galactic disc. MNRAS 346 (2), pp. L11 – L15 (English). External Links: Document Cited by: §7.2.
- Proper Motions and Orbits of Distant Local Group Dwarf Galaxies from a Combination of Gaia and Hubble Data. ApJ 971 (1), pp. 98. External Links: Document, 2312.09276 Cited by: §2.1, §2.2.
- The inefficiency of satellite accretion in forming extended star clusters. MNRAS 447 (1), pp. L40 – L44 (English). External Links: Document Cited by: §7.4.
- An Updated Small Magellanic Cloud and Magellanic Bridge Catalog of Star Clusters, Associations, and Related Objects. AJ 159 (3), pp. 82. External Links: Document, 1907.08642 Cited by: §7.2, §7.6.
- Astropy/photutils: 1.8.0. Zenodo. External Links: Document, Link Cited by: §5.1.
- PARSEC: stellar tracks and isochrones with the PAdova and TRieste Stellar Evolution Code. MNRAS 427 (1), pp. 127–145. External Links: Document, 1208.4498 Cited by: §5.2, §6, §7.1.
- Extragalactic Globular Clusters and Galaxy Formation. ARA&A 44 (1), pp. 193–267. External Links: Document, astro-ph/0602601 Cited by: §1.
- Stellar population synthesis at the resolution of 2003. MNRAS 344 (4), pp. 1000–1028. External Links: Document, astro-ph/0309134 Cited by: §5.2, §6, §7.2.
- X-ray spectral modelling of the AGN obscuring region in the CDFS: Bayesian model selection and catalogue. A&A 564, pp. A125. External Links: Document, 1402.0004 Cited by: §5.2.
- The Relationship between Infrared, Optical, and Ultraviolet Extinction. ApJ 345, pp. 245. External Links: Document Cited by: §5.2.
- ELVES II: globular clusters and nuclear star clusters of dwarf galaxies: the importance of environment. ApJ 927 (1), pp. 44. External Links: ISSN 0004-637X, Document, 2105.03440 Cited by: §7.6.
- Inferring the star formation histories of massive quiescent galaxies with BAGPIPES: evidence for multiple quenching mechanisms. MNRAS 480 (4), pp. 4379–4401. External Links: Document, 1712.04452 Cited by: §5.2, §6.
- DELVE 6: an ancient, ultra-faint star cluster on the outskirts of the magellanic clouds. ApJ 953 (2), pp. L21. External Links: ISSN 2041-8205, Document Cited by: footnote 9.
- Discovery of an Ultra-faint Stellar System near the Magellanic Clouds with the DECam Local Volume Exploration Survey. ApJ 910 (1), pp. 18. External Links: Document, 2009.08550 Cited by: footnote 9.
- Spectroscopy of Star Cluster Candidates and H II Regions in NGC 6822. AJ 120 (6), pp. 3088–3097. External Links: Document Cited by: §2.1, Table 7, §6.
- The Stellar Content of Nearby Star-forming Galaxies. III. Unravelling the Nature of the Diffuse Ultraviolet Light. ApJ 628 (1), pp. 210–230. External Links: Document, astro-ph/0505024 Cited by: §7.2.
- The Age, Mass, and Size Distributions of Star Clusters in M51. ApJ 824 (2), pp. 71. External Links: Document Cited by: §7.4.
- The Luminosity, Mass, and Age Distributions of Compact Star Clusters in M83 Based on Hubble Space Telescope/Wide Field Camera 3 Observations. ApJ 719 (1), pp. 966–978. External Links: Document, 1007.5237 Cited by: §5.1.
- PARSEC evolutionary tracks of massive stars up to 350 M⊙ at metallicities 0.0001 Z 0.04. MNRAS 452 (1), pp. 1068–1080. External Links: Document, 1506.01681 Cited by: §6, §7.1.
- DDO 216-a1: a central globular cluster in a low-luminosity transition-type galaxy. ApJ 837 (1), pp. 54. External Links: Document Cited by: §7.6.
- The Propagation of Uncertainties in Stellar Population Synthesis Modeling. I. The Relevance of Uncertain Aspects of Stellar Evolution and the Initial Mass Function to the Derived Physical Properties of Galaxies. ApJ 699 (1), pp. 486–506. External Links: Document, 0809.4261 Cited by: §5.2, §7.2.
- The Propagation of Uncertainties in Stellar Population Synthesis Modeling. III. Model Calibration, Comparison, and Evaluation. ApJ 712 (2), pp. 833–857. External Links: Document, 0911.3151 Cited by: §5.2, §7.2.
- Star cluster catalogues for the LEGUS dwarf galaxies. MNRAS 484 (4), pp. 4897–4919. External Links: Document, 1902.00082 Cited by: §5.1.
- The Formation of Giant Elliptical Galaxies and Their Globular Cluster Systems. ApJ 501 (2), pp. 554–570. External Links: Document, astro-ph/9804319 Cited by: §1.
- Deep Imaging of Eridanus II and Its Lone Star Cluster. ApJ 824 (1), pp. L14. External Links: Document, 1604.08590 Cited by: §1.
- Euclid: Early Release Observations – Programme overview and pipeline for compact- and diffuse-emission photometry. A&A 697, pp. A6. External Links: Document, 2405.13496 Cited by: §3.1, §3.1, §3.2, §3, §5.1.
- Evidence for Tidal Interaction and a Supergiant H I Shell in the Local Group Dwarf Galaxy NGC 6822. ApJ 537 (2), pp. L95–L98. External Links: Document, astro-ph/0005473 Cited by: §2.1.
- Satellite Dwarf Galaxies in a Hierarchical Universe: The Prevalence of Dwarf-Dwarf Major Mergers. ApJ 794 (2), pp. 115. External Links: Document, 1406.3344 Cited by: §1.
- Bright, relatively isolated star clusters in PHANGS-HST galaxies: Aperture corrections, quantitative morphologies, and comparison with synthetic stellar population models. MNRAS 510 (1), pp. 32–53. External Links: Document, 2110.10708 Cited by: §5.1.
- Evolved stars in the Local Group galaxies - II. AGB, RSG stars, and dust production in IC10. MNRAS 479 (4), pp. 5035–5048. External Links: Document, 1806.04160 Cited by: §2.2.
- A Carbon star approach to IC 10: Distance and correct size. A&A 424, pp. 125–132. External Links: Document Cited by: §2.2, §2.2.
- A Local Group Polar Ring Galaxy: NGC 6822. ApJ 636 (2), pp. L85–L88. External Links: Document, astro-ph/0510778 Cited by: §2.1.
- The ACS Survey of Galactic Globular Clusters. II. Stellar Evolution Tracks, Isochrones, Luminosity Functions, and Synthetic Horizontal-Branch Models. AJ 134 (1), pp. 376–390. External Links: Document, 0706.0847 Cited by: §1.
- Clearing the Hurdle: The Mass of Globular Cluster Systems as a Function of Host Galaxy Mass. ApJ 926 (2), pp. 162. External Links: Document, 2110.15376 Cited by: §7.6.
- The Recent Star Formation in NGC 6822: An Ultraviolet Study. ApJ 730 (2), pp. 88. External Links: Document, 1101.6051 Cited by: §2.1.
- A Universal Formation Mechanism for Open and Globular Clusters in Turbulent Gas. ApJ 480 (1), pp. 235–245. External Links: Document Cited by: §7.2.
- The Structure of Young Star Clusters in the Large Magellanic Cloud. ApJ 323, pp. 54. External Links: Document Cited by: §6.
- Euclid - ii. the vis instrument. A&A 697, pp. A2. External Links: Document, Link Cited by: §3.1.
- Euclid - iii. the nisp instrument. A&A 697, pp. A3. External Links: Document, Link Cited by: §3.1.
- Euclid - i. overview of the euclid mission. A&A 697, pp. A1. External Links: Document, Link Cited by: §1.
- Euclid preparation. I. The Euclid Wide Survey. A&A 662, pp. A112. External Links: Document, 2108.01201 Cited by: §3.1.
- Note: https://doi.org/10.57780/esa-qmocze3 Cited by: footnote 1.
- Dynamical Evolution of the Mass Function of Globular Star Clusters. ApJ 561 (2), pp. 751–765. External Links: Document, astro-ph/0107298 Cited by: §7.2.
- MULTINEST: an efficient and robust Bayesian inference tool for cosmology and particle physics. MNRAS 398 (4), pp. 1601–1614. External Links: Document, 0809.3437 Cited by: §5.2.
- Multimodal nested sampling: an efficient and robust alternative to Markov Chain Monte Carlo methods for astronomical data analyses. MNRAS 384 (2), pp. 449–463. External Links: Document, 0704.3704 Cited by: §5.2.
- PHANGS hubble space telescope treasury survey: globular cluster systems in 17 nearby spiral galaxies. AJ 167 (3), pp. 95. External Links: ISSN 0004-6256, Document, 2403.13908 Cited by: §7.6.
- Globular cluster formation and evolution in the context of cosmological galaxy assembly: open questions. Proceedings of the Royal Society of London Series A 474 (2210), pp. 20170616. External Links: Document, 1801.05818 Cited by: §1.
- Extending the globular cluster system-halo mass relation to the lowest galaxy masses. MNRAS 481 (4), pp. 5592–5605. External Links: Document, 1809.07831 Cited by: Figure 19, §7.6, §7.6, §7.6.
- Kinematics and simulations of the stellar stream in the halo of the Umbrella Galaxy. MNRAS 442 (4), pp. 3544–3564. External Links: Document, 1406.5511 Cited by: §1.
- Accounting for stochastic fluctuations when analysing the integrated light of star clusters. I. First systematics. A&A 521, pp. A22. External Links: Document, 1003.2334 Cited by: Appendix A, Appendix A.
- Distance and reddening of the Local Group dwarf irregular galaxy NGC 6822. A&A 548, pp. A129. External Links: Document, 1210.6753 Cited by: §2.1, §2.1, Table 1, §7.1, §7.2.
- A state-of-the-art analysis of the dwarf irregular galaxy NGC 6822. A&A 572, pp. A26. External Links: Document, 1409.5247 Cited by: §2.1.
- The Local Group Dwarf Irregular Galaxy NGC 6822.I.The Stellar Content. AJ 112, pp. 1928. External Links: Document Cited by: §2.1, §2.1.
- Nitrogen in Irregular Galaxies. ApJ 363, pp. 142. External Links: Document Cited by: §2.2.
- STEP survey - II. Structural analysis of 170 star clusters in the SMC. MNRAS 507 (3), pp. 3312–3330. External Links: Document, 2108.02791 Cited by: Figure 11, Figure 12, §7.2, §7.2, §7.4.
- Globular cluster systems in nearby dwarf galaxies - III. Formation efficiencies of old globular clusters. MNRAS 406 (3), pp. 1967–1984. External Links: Document, 1004.2039 Cited by: §7.6.
- Globular cluster systems in nearby dwarf galaxies - I. HST/ACS observations and dynamical properties of globular clusters at low environmental density. MNRAS 392 (2), pp. 879–893. External Links: Document, 0810.3660 Cited by: Figure 14, Figure 15, §7.3, §7.3, §7.4, §7.4.
- The red extended structure of IC 10, the nearest blue compact galaxy. MNRAS 454 (1), pp. 1000–1011. External Links: Document, 1509.05436 Cited by: §2.2, Table 1, Figure 13, §7.2.1.
- The Araucaria Project: An Accurate Distance to the Local Group Galaxy NGC 6822 from Near-Infrared Photometry of Cepheid Variables. ApJ 647 (2), pp. 1056–1064. External Links: Document, astro-ph/0605231 Cited by: §2.1.
- Age Determination of Six Intermediate-Age Small Magellanic Cloud Star Clusters with HST/ACS. AJ 136 (4), pp. 1703–1727. External Links: Document, 0807.3744 Cited by: §7.6.
- Hierarchical Stellar Structures in the Local Group Dwarf Galaxy NGC 6822. ApJ 725 (2), pp. 1717–1734. External Links: Document, 1010.1940 Cited by: footnote 4.
- The spatial relation between young star clusters and molecular clouds in M51 with LEGUS. MNRAS 483 (4), pp. 4707–4723. External Links: Document, 1812.06109 Cited by: §1.
- What is a globular cluster? an observational perspective. A&ARv 27 (1), pp. 8 – 136 (English). External Links: Document Cited by: §7.6.
- dustmaps: A Python interface for maps of interstellar dust. The Journal of Open Source Software 3 (26), pp. 695. External Links: Document Cited by: item 3.
- Rubin-Euclid Derived Data Products: Initial Recommendations. In Zenodo id. 5836022, Vol. 58, pp. 5836022. External Links: Document, 2201.03862 Cited by: §8.
- Array programming with NumPy. Nature 585 (7825), pp. 357–362. External Links: Document, Link Cited by: Data Availability.
- Globular clusters in galaxies beyond the local group. I. New cluster systems in selected northern ellipticals.. AJ 86, pp. 1627–1642. External Links: Document Cited by: §7.6.
- A Catalog of Parameters for Globular Clusters in the Milky Way. AJ 112, pp. 1487. External Links: Document Cited by: §7.4.
- Solo dwarfs II: the stellar structure of isolated Local Group dwarf galaxies. MNRAS 503 (1), pp. 176–199. External Links: Document, 2101.03189 Cited by: Table 1, Figure 13.
- Dusty Stellar Birth and Death in the Metal-poor Galaxy NGC 6822. ApJ 892 (2), pp. 91. External Links: Document, 2002.12437 Cited by: §2.1, §7.2.1.
- The structure and content of NGC 6822.. ApJS 33, pp. 69. External Links: Document Cited by: §2.1, §2.1, §4.2, footnote 2, footnote 4.
- The recent evolutionary history of the galaxies NGC 6822 and IC 1613.. ApJ 241, pp. 125–131. External Links: Document Cited by: §2.1.
- The H II Regions of NGC 6822. I. an Atlas of 157 H II Regions. PASP 100, pp. 917. External Links: Document Cited by: §2.1.
- NGC 6822, a remote stellar system.. ApJ 62, pp. 409–433. External Links: Document Cited by: §2.1, §4.2, footnote 2.
- Euclid: Early Release Observations – Deep anatomy of nearby galaxies. A&A 697, pp. A9. External Links: Document, 2405.13499 Cited by: Table 1, §3.1, §4.1, §7.2.1.
- The Stellar Population and Star Clusters in the Unusual Local Group Galaxy IC 10. ApJ 559 (1), pp. 225–242. External Links: Document, astro-ph/0105456 Cited by: §2.2, §4.2, footnote 3, footnote 8.
- N-body models of extended star clusters. MNRAS 408 (4), pp. 2353–2363. External Links: Document, 1008.4991 Cited by: §7.4.
- Three newly discovered globular clusters in NGC 6822. MNRAS 429 (2), pp. 1039–1044. External Links: Document, 1211.3896 Cited by: §2.1, §4.2, Table 6, §7.1, §7.2, §7.4, footnote 2.
- The outer halo globular cluster system of M31 - I. The final PAndAS catalogue. MNRAS 442 (3), pp. 2165–2187. External Links: Document, 1404.5807 Cited by: §1, §1, §4.1, §7.3.
- Globular clusters in the outer halo of M31: the survey. MNRAS 385 (4), pp. 1989–1997. External Links: Document, 0801.0002 Cited by: §4.1.
- A new population of extended, luminous star clusters in the halo of M31. MNRAS 360 (3), pp. 1007–1012. External Links: Document, astro-ph/0412223 Cited by: §1, §7.4.
- Extended Star Clusters in the Remote Halo of the Intriguing Dwarf Galaxy NGC 6822. ApJ 738 (1), pp. 58. External Links: Document, 1106.2878 Cited by: §2.1, §4.2, Table 6, §7.4, footnote 2.
- Spectroscopic Study of Extended Star Clusters in Dwarf Galaxy NGC 6822. ApJ 783 (1), pp. 49. External Links: Document, 1401.2492 Cited by: §2.1, Table 7, §6, §7.1, §7.6.
- Discovery of the Most Isolated Globular Cluster in the Local Universe. ApJ 751 (1), pp. L19. External Links: Document, 1205.3559 Cited by: §1, §7.4.
- The 2MASS Large Galaxy Atlas. AJ 125 (2), pp. 525–554. External Links: Document Cited by: Figure 13.
- Panchromatic Hubble Andromeda Treasury. XVIII. The High-mass Truncation of the Star Cluster Mass Function. ApJ 839 (2), pp. 78. External Links: Document, 1703.10312 Cited by: §1, §7.2.
- PHAT Stellar Cluster Survey. II. Andromeda Project Cluster Catalog. ApJ 802 (2), pp. 127. External Links: Document, 1501.04966 Cited by: §1, §4.1.
- The Panchromatic Hubble Andromeda Treasury: Triangulum Extended Region (PHATTER). IV. Star Cluster Catalog. ApJ 938 (1), pp. 81. External Links: Document, 2208.11760 Cited by: Appendix A, §1.
- Gas-rich, Field Ultra-diffuse Galaxies Host Few Globular Clusters. ApJ 942 (1), pp. L5. External Links: Document, 2211.00651 Cited by: §7.6.
- Photometric distances to the nearby galaxies IC 10, IC 342, and UGCA 86, visible through the Milky Way.. A&AS 100, pp. 227–235. Cited by: §2.2, §4.2, footnote 3.
- Discovery of Globular Cluster Candidates in the Dwarf Irregular Galaxy IC 2574 Using HST/ACS Imaging. MNRAS 530 (4), pp. 4936–4949. External Links: Document, 2402.16955 Cited by: §7.6.
- Star Formation History of the Local Group Dwarf Irregular Galaxy, NGC 6822. arXiv e-prints, pp. arXiv:2412.05646. External Links: Document, 2412.05646 Cited by: §2.1.
- Reddening and Distance of the Local Group Starburst Galaxy IC 10. ApJ 703 (1), pp. 816–828. External Links: Document, 0907.4844 Cited by: §2.2.
- The structure of star clusters. III. Some simple dynamical models. AJ 71, pp. 64. External Links: Document Cited by: §6.
- The Universal Stellar Mass-Stellar Metallicity Relation for Dwarf Galaxies. ApJ 779 (2), pp. 102. External Links: Document, 1310.0814 Cited by: §2.1.
- Newly Identified Star Clusters in NGC 6822, and the Age Distribution of Its Cluster System. PASP 116 (820), pp. 497–505. External Links: Document Cited by: §2.1, §4.2, Table 6, footnote 2.
- On the variation of the initial mass function. MNRAS 322 (2), pp. 231–246. External Links: Document, astro-ph/0009005 Cited by: §5.2, item 1.
- The minimum metallicity of globular clusters and its physical origin - implications for the galaxy mass-metallicity relation and observations of proto-globular clusters at high redshift. MNRAS 486 (1), pp. L20–L25. External Links: Document, 1904.09987 Cited by: §7.6.
- Star Clusters Across Cosmic Time. MNRAS 57, pp. 227–303. External Links: Document, 1812.01615 Cited by: §1.
- The formation, evolution and disruption of star clusters with improved gravitational dynamics in simulated dwarf galaxies. MNRAS. External Links: Document Cited by: §1, §7.2.
- Astrometry.net: Blind Astrometric Calibration of Arbitrary Astronomical Images. AJ 139 (5), pp. 1782–1800. External Links: Document, 0910.2233 Cited by: §3.2.
- The chemical composition of globular clusters in the Local Group. A&A 660, pp. A88. External Links: Document, 2112.00081 Cited by: §2.1, Table 7, §6, §7.1, §7.2, §7.6.
[…]